OpenStudy (anonymous):

which is the equation of a line that has an x-intercept of -3 and is perpendicular to the line 2x-y=1? How do you solve it?

OpenStudy (anonymous):

the given line slope is 2 , our line slope is perp on it then our line slope is -2

OpenStudy (anonymous):

our line has x intercept at -3 then point (-3,0) lyes on it

OpenStudy (anonymous):

what would the equation be?

OpenStudy (anonymous):

equation should be y / (x+3) = -2

OpenStudy (anonymous):

y = -2x-6

OpenStudy (anonymous):

wait so how exactly did you get that?

OpenStudy (anonymous):

i already said how i got it :)

OpenStudy (anonymous):

fine

OpenStudy (anonymous):

bnb is incorrect. if the slope of a given line is m, then the perpendicular to that line will have a slope -1/m

OpenStudy (anonymous):

so the equation will be y/(x+3)=-1/2 and dat'll giv us 2y+x+3=0.

OpenStudy (anonymous):

x+2y+6=0

OpenStudy (anonymous):

ops my bad , thanx brackett

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

ops my bad , thanx chutzpah!